# NAAQA: A Neural Architecture for Acoustic Question Answering

### v2

@@ 6.1.2 Coordinate Maps
Coordinate maps can be inserted before any convolution operation (Figure 2). We therefore analyzed the impact of the placement of Time and Frequency coordinate maps at different depths in the network. All possible locations were evaluated via grid-search. For each location, we inserted either a Time coordinate map, a Frequency coordinate map or both. Results are detailed in Table V. Time coordinate maps have the biggest impact on performance, especially when inserted in the first convolution after the feature extractor or in the Resblocks. This could be because the fusion of the textual and acoustic features, and therefore most of the reasoning, is performed in the Resblocks. The network might be using the additional localization information to inform the modulation of the convolutional feature maps based on the context of the question. Surprisingly, the Frequency coordinate maps have a minimal impact on performance. We further compare the impact of Time versus Frequency coordinate maps in Supplementary B.
@@ 6.1.3 Complexity Optimization
Figure 4: Test accuracy by question type and the number of relation in the question for Optimized NAAQA Parallel. The overall accuracy for this configuration is 79.1%. The presence of before or after in a question constitutes a temporal relation. The accuracy is N/A for relative position and count compare since these types of question do no include relations. The hyper-parameters used are described at the end of Section 6.1.3
As described in Section 4.2.3, we optimized the most important hyper-parameters ($G,J,M$) in the NAAQA model in order to reduce its complexity. The baseline Visual FiLM Resnet101 configuration comprises 6.71 M parameters and achieves only 62.3%. NAAQA Parallel comprises 5.61 M parameters and performs significantly better with 78.5%. With this model as starting point, we performed an ablation study to find which hyper-parameters can be reduced without impacting accuracy. The Optimized NAAQA Parallel configuration is the best trade-off between model complexity and performance. It comprises 1.68 M parameters and achieves the best overall accuracy with 79.5%. The most notable complexity reduction comes from the reduction of the number of GRU units $G$. Reducing $G$ from $4096$ to $512$ increased accuracy while reducing the number of parameters by a factor of 3 (6.61 M vs 1.68 M). The Optimized NAAQA Parallel is composed of a Parallel extractor with $K=3$ blocks and $P=64$, $G=512$ GRU units, $J=4$ Resblocks with $M=128$ filters, a classifier module with $C=512$ filters and $H=1024$ units. Results for this configuration can be found in Table III and Figure 4. Further results related to the ablation study can be found in Supplementary H.
@@ 6.2 Summary of Results on CLEAR2
NAAQA performs well on the CLEAR2 AQA task with 79.5% overall accuracy. It does however struggle with certain types of question as shown in Table III and Figure 4. When asked to count the number of sounds with specific attributes, NAAQA reaches only 53.8% accuracy. This limitation is more severe if the question is to count the different instruments playing in a given part of the scene (50.4%). It attains slightly higher accuracy when asked to compare the number of instances of acoustic objects (more, fewer or equal number) with specific attributes (60.6%). In contrast, the network can successfully recognize individual instruments in the scene (81.7%). This suggests, that the problem lies in the logical complexity of the question rather than in the pattern matching from the acoustic scene. As an example, the question (count instrument): ”How many different instruments are playing after the third cello playing a C# note?” requires to first identify the cello playing the C# note, then identify all acoustic objects that are playing after this sound, determine which instruments are of the same family and finally count the number of different families. The model struggles when it must focus on a large number of acoustic objects which explain the low accuracy for this type of question.
A similar argument could explain why models also have difficulties with questions related to the relative position of the instruments (58.0%). For example, to answer the question ”Among the flute sounds, which one plays an F note?”, the model must find all the flutes playing in the scene, determine which one play an F note, count the number of flute playing before and translate the count to a position [^4] This is only one possible strategy to answer the question. There may be other ways.. This also requires the network to focus on multiple objects.
@@
6.3 Evaluation of DAQA
In order to compare our results to those of [undefaa] [undefaa], we evaluated our models on a version of the DAQA data set. As mentioned in Section 3.2, we were not able to reproduce the original DAQA dataset which mean that the results presented in this section are not fully comparable with [undefaa]. We report results for different configurations of NAAQA on our modified DAQA in Table IV.
@@ 6.3.1 NAAQA on DAQA
The models explored in this section are considerably smaller in terms of number of parameters than the ones explored in [undefaa]. The smallest model they evaluated had 5.49M parameters, the biggest model had 21.33M parameters and the best performing model had 13.20M parameters. The Optimized NAAQA 2D Conv model only has 1.68 M parameters and reaches an accuracy of 58.3% on DAQA. The Optimized NAAQA Parallel has the same number of parameters and performs slightly better with and accuracy of 60.4%. When we analyzed both of these models on CLEAR2 dataset in Section 6.1.1, we found a much smaller difference between the performance of the NAAQA 2D Conv and the NAAQA Parallel. This difference in the observations suggests that the parallel extractor is more useful in the context of complex acoustic sounds (DAQA) than with sustained musical note (CLEAR2).
Even though these results are not 100% comparable with [undefaa] because of the difference in the dataset composition, we want to emphasize that Optimized NAAQA Parallel reaches a somewhat similar accuracy than the smallest FiLM in [undefaa] (60.4% vs 64.3%) with significantly smaller number of parameters (1.68 M vs 5.49 M).
@@ 6.3.2 NAAQA with a MALiMo module on DAQA
In Section 6.1.4, we found that adding a MALiMo controller to our NAAQA models did not improve the accuracy on CLEAR2. Results in Table IV show that the MALiMo controller have a significant impact on performance when the model is evaluated on DAQA dataset. We see an increase of almost 4% when using Optimized NAAQA Parallel + MALiMo ctrl compared to Optimized NAAQA Parallel alone. These results are consistent with [undefaa] findings and with the hypothesis that MALiMo increases performance when working with complex sounds.
The Optimized NAAQA Parallel + MALiMo ctrl configuration performs about the same as the smallest MALiMo model evaluated in [undefaa] (64.3% vs 65.1%) with significantly less parameters (2.78 M vs 8.91 M).
@@ 7 Conclusions
Acoustic Question Answering (AQA) is a newly emerging task in the area of machine learning research. As performance is strongly dependent on the acoustical environments and types of questions, it is important to understand the relationship between the application and the chosen neural architecture. We propose a benchmark for AQA based on musical sounds (CLEAR2) and a neural architecture that is tailored to interpreting acoustic scenes (NAAQA). NAAQA introduces a number of modifications to a FiLM based architecture in order to optimize acoustic scenes analysis. These includes several strategies for neural feature extraction, an ablation study of the hyper-parameters and the optimization of coordinate maps for this task. We confirm that FiLM layers are very effective to modulate activation maps in the AQA application. We are able to optimize our NAAQA neural network so to obtain competitive results with a fraction of the model complexity. These results are confirmed on a different AQA task (DAQA) comprising more complex sounds with the addition of a MALiMo controller in the model. We release all code openly in the hope that these resources may foster increased research activity in solving the AQA task.
@@ Acknowledgment
We would like to acknowledge the NVIDIA Corporation for donating a number of GPUs. Part of this research was funded or supported by the CHIST-ERA IGLU project, the CRSNG, the Michael-Smith scholarships the University of Sherbrooke and NTNU.
Supplementary Material
Supplementary A Detailed training settings
All models were trained for a maximum of 40 epochs or until the validation loss stops decreasing for 6 consecutive epochs. The Adam optimizer [undefaav] with a learning rate of $3\text{\times}{10}^{-4}$ was used to optimize the cross entropy loss. To reduce the likelihood of the model getting trapped in a local minima, the learning rate is automatically reduced by a 0.1 factor if the validation loss stops decreasing for 3 consecutive epochs. L2 regularization with a weight decay of $5\text{\times}{10}^{-6}$, dropout [undefaaw] with a drop probability of 0.25 and batch normalization [undefaap] are used to regularize the learning process. A batch size of 128 is used for all experiments. The following 5 random seeds were used to run all experiments : 876944 189369 682421 175326 427438
Supplementary B Time and Frequency coordinate maps impact
Accuracy by question type on the test set (%)
Coordinate Absolute Global Relative Count Count
Maps Overall Instrument Note Brightness Loudness Exist Position Position Position Count Comparison Instruments
Time only 79.0 ±0.43 80.8 ±1.22 73.0 ±0.84 90.8 ±0.52 86.8 ±0.49 80.2 ±0.75 80.2 ±0.89 89.0 ±0.60 53.1 ±0.90 52.1 ±0.70 61.1 ±0.56 49.3 ±0.47
Time & Freq 78.7 ±0.70 80.3 ±0.92 72.5 ±0.82 90.8 ±0.83 87.0 ±1.02 80.6 ±0.47 79.9 ±1.25 88.9 ±0.58 54.8 ±1.21 52.9 ±0.42 59.1 ±2.46 46.0 ±3.14
None 63.8 ±0.70 73.3 ±0.72 61.8 ±0.62 87.6 ±0.66 84.6 ±0.96 77.1 ±0.56 41.0 ±2.19 69.8 ±1.33 49.0 ±2.02 45.9 ±0.80 59.3 ±0.41 44.6 ±1.71
Freq only 61.7 ±1.16 72.7 ±0.88 60.4 ±0.89 87.4 ±0.62 83.4 ±0.41 75.6 ±0.71 36.8 ±3.80 67.2 ±1.58 47.6 ±1.86 44.2 ±1.58 57.0 ±2.80 43.5 ±4.09
TABLE VI: Comparison of Time and Frequency coordinate maps. Coordinate maps are inserted in the Resblocks. Test accuracy is reported by question type (Examples for each question type are available in Table I).
The impact of the kind of coordinate maps on each question types is presented in Table VI. To facilitate comparison, all coordinate maps are inserted at the same location (in the first convolution after the feature extractor). When the network does not include any coordinate maps it reaches only 63.8%. Inserting a Frequency coordinate map decreases accuracy to 61.7%. As hypothesized, Time coordinate maps significantly increase the overall accuracy with 79%. The accuracy increase for all question types is most notable with questions related to Absolute position where the accuracy goes from 41.0% to 80.2%. This is a similar result to [undefaaj] where coordinate maps were proved to help convolutional networks locate objects in visual scenes. In the context of a spectro-temporal representation, this ability translates to enhanced temporal localization which naturally helps to answer questions related to position. Questions related to Notes also receive a significant boost from Time coordinate maps going from 61.8% to 73%. The fact that Frequency coordinate maps decrease performance on questions related to Notes is rather surprising since notes are mostly defined by the frequency content at a given point in time. Furthermore, combining both Time and Frequency coordinate maps slightly reduces the overall performance when compared to using only Time coordinate maps which is in line with previous results.
Supplementary C Results showing Effects of stride size on both CLEAR2 and DAQA’
Model Nb Parameters Train acc Val acc Test acc Trainig time
On CLEAR2 dataset
Optimized NAAQA Parallel- Long stride 1.68 M 95.2 ±0.43 90.5 ±0.28 79.1 ±0.32 0 days 02:39:51
Optimized NAAQA Parallel- Short stride 1.68 M 93.1 ±0.48 89.5 ±0.38 77.9 ±0.68 0 days 07:00:01
On DAQA dataset
Optimized NAAQA Parallel- Short Stride 1.68 M 68.0 ±0.22 64.6 ±0.34 65.3 ±0.37 3 days 17:27:45
Optimized NAAQA Parallel- Long Stride 1.68 M 67.3 ±1.38 60.2 ±0.24 60.3 ±0.20 0 days 07:14:17
TABLE VII: Short stride and Long Stride comparison The optimized NAAQA on CLEAR2 and DAQA’. The rows are ordered by test accuracy.
Supplementary D Discussion on results with Resnet
Multiple feature extractors were evaluated in Section 6.1.1. One of them is the Resnet-101 pre-trained on ImageNet (Visual dataset). This extractor achieves the lowest of all tested extractors with 62.3%. However, it performs quite well with questions related to brightness and loudness with 83.9% and 81.9% respectively. One possible explanation is that both brightness and loudness are directly related to the amplitude of the “pixels” along the frequency axis which may be efficiently encoded by the Resnet101 model. The baseline does not perform as well with questions related to absolute position which is somewhat surprising since these questions correspond to a similar pattern matching problem on the time axis instead of the frequency axis. The baseline model also has difficulties with questions related to notes (37.8%). This might be explained by the fact that a note can be identified by its fundamental frequency and harmonics which can be far apart on the frequency axis. Visual models are trained to recognize localized features and therefore struggle to recognize notes. This is a feature that is typical of acoustic signals.
It does however perform better with global position questions. This is logical because this type of question refers to an approximate position (beginning, middle, end) which constrain the number of possible answers to 3 instead of 15.
Supplementary E Variability of the input size
Batch normalization has been proven to help regularize the training of neural networks [undefaap]. However, using batch normalization requires all inputs in a batch to have the same dimensions in order to compute the element-wise means and standard deviations. Acoustic signals in the CLEAR dataset have variable duration. To work around this constraint, we padded the spectrograms with zeros so that they all have the same dimensions. Another solution is to resize all spectrograms to a fixed size using bilinear interpolation as it is commonly done in the visual domain. From a purely acoustic point of view, we hypothesized that padding spectrograms would preserve the relative time axis between all scenes and would yield better results than resizing. We revisited this hypothesis by training our best architecture on resized spectrograms. Surprisingly, the network performed better by about 1 percentage points when trained on the resized spectrograms. We do not have a definitive answer as to why we observe this behavior. It might be due to the fact that padded spectrograms contain less information since a portion of them is filled with zeros. The network could also be using the different time resolutions in the resized spectrograms as a cue to cheat the reasoning process. However, this could also happen when the scenes are padded since the padded section is unique to each scene. We analyzed the network activations and did not find that the padded section was overly activated. We leave it up to future work to further analyze this.
Supplementary F Importance of text versus audio modality
To validate the fact that NAAQA makes good use of both input modalities (spectrogram and text), the network was trained using only one modality. To do so, either all values of the input spectrograms were set to 1 or all questions were set to the unknown token during training. No more than 48% of accuracy on the test set was achieved for either of these scenarios which indicates that the network uses information from both modalities to achieve high performance and that the dataset cannot be solved using only one modality with this architecture. This is however much higher than the probability of answering the correct answer by chance ($\frac{1}{57}=1.75\%$), or by majority vote (7.5%) introduced in Section 3.1. Although we have carefully designed the data set in order to avoid bias, some form of bias is intrinsic in QA problems. For example, an acoustic sequence containing only $n$ elementary sounds automatically disables all counting answers above $n$ regardless which question is asked. Similarly, asking a yes/no question limits the number of possible answers to two, regardless of the acoustic scene.
Supplementary G DAQA’ missing sounds
Here is a list of all the missing elementary sounds that were originally in DAQA and that are not in the DAQA’ dataset (because of not being anymore accessible on youtube) :
b000_20, b000_3, b000_4, b000_5, b000_6, b001_11, c001_13, c001_14, c001_15, c001_16, c001_17, c001_18, c001_19, c001_20, c003_9, c004_10, d001_1, d001_9, d002_19, f001_19, h001_10, h001_11, h001_12, h001_13, h001_14, h001_15, h001_19, h001_2, h001_3, h001_4, h001_5, h001_6, h001_7, h001_8, h001_9, h002_10, h002_11, h002_12, h002_8, h002_9, t000_14, t000_15
Supplementary H Network Optimization
Gru (G) Overall Accuracy (%) Nb of
Units Train Val Test Parameters
512 95.5 ±0.28 90.1 ±0.37 79.6 ±0.30 1.94 M
256 94.3 ±0.51 90.5 ±0.30 79.5 ±0.05 1.68 M
1,024 94.2 ±0.31 89.7 ±0.18 79.0 ±0.14 2.46 M
4,096 94.8 ±0.25 89.9 ±0.25 78.9 ±0.23 5.61 M
2,048 95.8 ±0.14 90.0 ±0.05 78.7 ±0.04 3.51 M
TABLE VIII: Impact of the number of GRU text-processing units $G$. The interleaved extractor is used with hyper-parameters from the initial configuration described in Section 5.3 except for $G$. The number of GRU unit has a strong impact on the number of parameters. Setting $G=512$ actually increases accuracy while reducing the number of parameters by a factor of $\sim$2.8. The rows are ordered by test accuracy.
Nb (J) Nb Filter (M) Overall Accuracy (%) Nb of
Resblock in Resblock Train val Test Parameters
4 128 94.9 ±0.91 89.9 ±0.39 79.1 ±0.56 1.94 M
3 128 95.0 ±0.67 89.4 ±0.42 78.4 ±0.25 1.64 M
2 128 93.6 ±0.40 86.9 ±0.61 77.3 ±0.88 1.35 M
4 64 92.8 ±0.95 87.6 ±0.98 77.0 ±0.66 1.11 M
3 64 91.8 ±1.36 86.2 ±0.77 76.4 ±0.65 1.01 M
2 64 90.8 ±0.75 84.5 ±0.64 75.0 ±0.63 0.90 M
4 32 88.1 ±1.35 83.1 ±1.08 73.6 ±1.18 0.82 M
3 32 86.3 ±1.39 80.6 ±1.54 71.5 ±1.54 0.78 M
2 32 84.4 ±1.36 76.8 ±2.02 68.4 ±1.59 0.74 M
1 128 89.0 ±1.80 75.3 ±2.65 67.2 ±2.43 1.05 M
1 64 84.3 ±4.61 69.6 ±4.61 62.1 ±3.41 0.79 M
1 32 67.0 ±3.58 57.3 ±2.07 52.9 ±1.50 0.69 M
TABLE IX: Impact of reducing the number of Resblock $J$ and the number of filters $M$ in each Resblocks. The interleaved extractor is used with hyper-parameters from the initial configuration described in Section 5.3 except for $J$, $M$, $C=256$, $H=1024$ and $G=512$. With $J=4$ and $M=64$, the network reaches great accuracy with only 0.67M parameters (vs 5.61M for the initial configuration). The accuracy drops when $J<3$ and $M=32$. The configuration $J=3$ and $M=128$ get stuck in local minimums often which explain the large STD. The rows are ordered by test accuracy
Tables VIII and IX report the intermediate results of the NAAQA network complexity study on CLEAR2.

### v3

@@ 6.1.2 Coordinate Maps
Coordinate maps can be inserted before any convolution operation (Figure 2). We therefore analyzed the impact of the placement of Time and Frequency coordinate maps at different depths in the network. All possible locations were evaluated via grid-search. For each location, we inserted either a Time coordinate map, a Frequency coordinate map or both. Results are detailed in Table V. Time coordinate maps have the biggest impact on performance, especially when inserted in the first convolution after the feature extractor or in the Resblocks. This could be because the fusion of the textual and acoustic features, and therefore most of the reasoning, is performed in the Resblocks. The network might be using the additional localization information to inform the modulation of the convolutional feature maps based on the context of the question. Surprisingly, the Frequency coordinate maps have a minimal impact on performance. We further compare the impact of Time versus Frequency coordinate maps in Supplementary Materials.
@@ 6.1.3 Complexity Optimization
Figure 4: Test accuracy by question type and the number of relation in the question for Optimized NAAQA Parallel. The overall accuracy for this configuration is 79.1%. The presence of before or after in a question constitutes a temporal relation. The accuracy is N/A for relative position and count compare since these types of question do no include relations. The hyper-parameters are described in the end of Section 6.1.3
As described in Section 4.2.3, we optimized the most important hyper-parameters ($G,J,M$) in the NAAQA model to reduce its complexity. The baseline Visual FiLM Resnet101 configuration comprises 6.71 M parameters and achieves only 62.3%. NAAQA Parallel comprises 5.61 M parameters and performs significantly better with 78.5%. With this model as a starting point, we performed an ablation study to find which hyper-parameters can be reduced without impacting accuracy. The Optimized NAAQA Parallel configuration is the best trade-off between model complexity and performance. It comprises 1.68 M parameters and achieves the best overall accuracy with 79.5%. The most notable complexity reduction comes from the reduction of the number of GRU units $G$. Reducing $G$ from $4096$ to $512$ increased accuracy while reducing the number of parameters by a factor of 3 (6.61 M vs 1.68 M). The Optimized NAAQA Parallel is composed of a Parallel extractor with $K=3$ blocks and $P=64$, $G=512$ GRU units, $J=4$ Resblocks with $M=128$ filters, a classifier module with $C=512$ filters and $H=1024$ units. Results for this configuration can be found in Table III and Figure 4. Further results related to the ablation study can be found in Supplementary Materials.
@@ 6.2 Summary of Results on CLEAR2
NAAQA performs well on the CLEAR2 AQA task with 79.5% overall accuracy. It does however struggle with certain types of question as shown in Table III and Figure 4. When asked to count the number of sounds with specific attributes, NAAQA reaches only 53.8% accuracy. This limitation is more severe if the question is to count the different instruments playing in a given part of the scene (50.4%). It attains slightly higher accuracy when asked to compare the number of instances of acoustic objects (more, fewer or equal number) with specific attributes (60.6%). In contrast, the network can successfully recognize individual instruments in the scene (81.7%). This suggests, that the problem lies in the logical complexity of the question rather than in the pattern matching from the acoustic scene. As an example, the question (count instrument): ”How many different instruments are playing after the third cello playing a C# note?” requires to first identify the cello playing the C# note, then identify all acoustic objects that are playing after this sound, determine which instruments are of the same family and finally count the number of different families. The model struggles when it must focus on a large number of acoustic objects which explains the low accuracy for this type of question.
A similar argument could explain why models also have difficulties with questions related to the relative position of the instruments (58.0%). For example, to answer the question ”Among the flute sounds, which one plays an F note?”, the model must find all flutes playing in the scene, determine which one plays an F note, counts the number of flute playing before and translates the count to a position  [^7] This is one possible strategy to answer the question. There may be other ways.. This also requires the network to focus on multiple objects.
@@
6.3 Evaluation on DAQA
To compare our results to those of [undefab] [undefab], we evaluated our models on a version of the DAQA data set. As mentioned in Section 3.2, we were not able to reproduce the original DAQA dataset which means that results presented in this section are not fully comparable with [undefab]. Results for different configurations of NAAQA tested on our modified DAQA are reported in Table IV.
@@ 6.3.1 NAAQA on DAQA
The models explored in this section matches the performance of previous efforts [undefab]. The smallest model they evaluated had 5.49M parameters, the biggest model had 21.33M parameters and the best performing model had 13.20M parameters. The Optimized NAAQA 2D Conv model only has 1.68 M parameters and reaches an accuracy of 58.3% on DAQA. The Optimized NAAQA Parallel has the same number of parameters and performs slightly better with and accuracy of 60.4%. When we analyzed both of these models on CLEAR2 dataset in Section 6.1.1, we found a much smaller difference between the performance of the NAAQA 2D Conv and the NAAQA Parallel. This difference suggests that the parallel extractor is more effective in the context of complex acoustic sounds (DAQA) than with sustained musical notes (CLEAR2).
Even though these results are not 100% comparable with [undefab] because of the difference in the dataset composition, we want to emphasize that Optimized NAAQA Parallel reaches a somewhat similar accuracy than the smallest FiLM in [undefab] (60.4% vs 64.3%) with significantly smaller number of parameters (1.68 M vs 5.49 M).
@@ 6.3.2 NAAQA with a MALiMo module on DAQA
In Section 6.1.4, we found that adding a MALiMo controller to our NAAQA models did not improve the accuracy on CLEAR2. On the other hand, the MALiMo controller has a significant positive impact when the model is evaluated on DAQA dataset (Table IV). We see an increase of almost 4% when using Optimized NAAQA Parallel + MALiMo ctrl compared to Optimized NAAQA Parallel alone. These results are consistent with [undefab] findings and with the hypothesis that MALiMo increases performance when working with complex sounds.
The Optimized NAAQA Parallel + MALiMo ctrl configuration performs about the same as the smallest MALiMo model evaluated in [undefab] (64.3% vs 65.1%) with significantly fewer parameters (2.78 M vs 8.91 M).
@@ 7 Conclusions
Acoustic Question Answering (AQA) is a newly emerging task in the area of machine learning research. As performance is strongly dependent on the acoustical environments and types of questions, it is important to understand the relationship between the application and the chosen neural architecture. We propose a benchmark for AQA based on musical sounds (CLEAR2) and a neural architecture that is tailored to interpreting acoustic scenes (NAAQA). NAAQA introduces a number of modifications to a FiLM based architecture to optimize acoustic scenes analysis. These includes several strategies for neural feature extraction, an ablation study of the hyper-parameters and the optimization of coordinate maps. We confirm that FiLM layers are very effective to modulate activation maps in the AQA application. We are able to optimize our NAAQA neural network so to obtain competitive results with a fraction of the model complexity. These results are confirmed on a different AQA task (DAQA) comprising more complex sounds with the addition of a MALiMo controller in the model. We release all code openly in the hope that these resources may foster increased research activity in solving the AQA task.
@@ Acknowledgment
We would like to thank the reviewers for their constructive comments that helped us improve the paper. The NVIDIA Corporation for the donation of GPUs. Part of this research was funded or supported by the CHIST-ERA IGLU project, the CRSNG, the Michael-Smith scholarships, the Universities of Sherbrooke and NTNU.